OpenStudy (anonymous):

Help!!!!! Write an equation of the line that passes through the point (1,2) and has the slope, m = 2. I got Y=2X but it wasn't right.

OpenStudy (anonymous):

You need to use the point slope formula. \[y-y _{1}=m(x-x _{1})\] Where the point given represents \[(x _{1}, y _{1})\] and m is slope.

OpenStudy (anonymous):

I don't get it

OpenStudy (anonymous):

Do you know how to use formulas?

OpenStudy (anonymous):

I'm not good

OpenStudy (anonymous):

Ok, so you want to take what you have and decide what it will represent in your formula so that you can put it into your formula. So the point given is going to be the \[(x _{1}, y _{1})\] What are those two values?

OpenStudy (anonymous):

I don't know ? /:

OpenStudy (anonymous):

What is the point that is given to you in the problem?

OpenStudy (anonymous):

to use the m for the slope or somthing

OpenStudy (anonymous):

Points are given to you as two numbers inside of parenthesis with a comma separating them. Can you find that in the problem?

jimthompson5910 (jim_thompson5910):

The answer of y = 2x is correct. It has the slope of 2 and it goes through (1,2) y = 2x 2 = 2*1 ... plug in (x,y) = (1,2) 2 = 2 So this shows that y = 2x goes through (1,2)

jimthompson5910 (jim_thompson5910):

So you might have to write the answer in standard form or point slope form

OpenStudy (anonymous):

Well what I had put was not right

jimthompson5910 (jim_thompson5910):

So try 2x - y = 0 which is what you'd get if you convert to standard form

OpenStudy (anonymous):

I put 2x - y = 0 as the answer?

jimthompson5910 (jim_thompson5910):

yeah try that

OpenStudy (anonymous):

Thank you

jimthompson5910 (jim_thompson5910):

so that worked?

OpenStudy (anonymous):

I have other answers I have to redo , and then I can re submit my test :)

jimthompson5910 (jim_thompson5910):

ok glad it worked, i wish they explicitly stated they wanted the answer in standard form, but oh well

OpenStudy (anonymous):

haha yeah I know

OpenStudy (anonymous):

Think some. o.o